Two small balls A and B each of mass m are joined rigidly to the ends of a light rod of length L. The system translates on a frictionless horizontal surface with a velocity v in a direction perpendicular to the rod. A particle P of mass m kept at rest on the surface sticks to the ball A as the ball collides with it. Find (a) the linear speeds of the balls A and B after the colloision (b) the velocity of the centre of mass C of the system A+B+P and (c) the angular speed of the system about C after the colloision.


Grade:upto college level

1 Answers

Aman Sengar
7 years ago
Dear student, The linear speeds of balls can be calculated by conservation of linear momentum I.e.mv=2mv`V`=v/2 this is the velocity of partial a And since there is no external force acting on ball B at that instant hence velocity will remain unchanged I.e. v.Velocity of COM can be calculated by (M1v1+m2v2......)/M1+M1+......Hence Vcm= (mv+2mv/2)/3m =2v/3.And for last part we can use W=v (perpendicular)/rW= (v-v/2)/l = v/2lI think this might help you.